OpenStudy (anonymous):

How do I get started here? Help please. 14.72-21.58x=18/2.11x + 2.4

OpenStudy (anonymous):

Convert the decimal numbers to integer fractions.\[\frac{1472}{100}-\frac{2158 x}{100}=\frac{18 x}{\frac{211}{100}}+\frac{24}{10} \]Solve the above for x. The result is an exact symbolic solution. \[x=\frac{11816}{28879} \]x to 20 decimal digits.\[0.40915544167041795076 \]

OpenStudy (anonymous):

i am getting x =18816/28879 =0.4091554417

OpenStudy (anonymous):

0.40915544167 rounds to 0.4091554417